SPECTROSCOPY OF NbO: CHARACTERIZATION OF THE DOUBLET MANIFOLD \underline{O. LAUNILA}, U. WAHLGREN, B. SCHIMMELPFENNIG and A. G. TAKLIF, Department of Physics, Stockholm University, Box 6730 S-11385 Stockholm, Sweden.